# The Born rule – 100 years ago and today

### v1

@@
Abstract. Details of the contents, and formulations of the Born rule changed considerably from its inception by Born in 1926 to the present day. Based to a large extent on little known results from the recent books ’Coherent Quantum Physics’ by Neumaier and ’Algebraic Quantum Physics’, Vol. 1 by Neumaier and Westra, this paper traces the early history 100 years ago, its generalization (essential for today’s quantum optics and quantum information theory) to POVMs 50 year ago, and a modern derivation from an intuitive definition of the notion of a quantum detector. Also discussed is the extent to which the various forms of the Born rule have, like any other statement in physics, a restricted domain of validity, which leads to problems when applied outside this domain.
@@ 1 Ten formulations of the Born rule
For example, we read in Jeans [28, p.91] (1924): ”Obviously quantum-dynamics must be intimately concerned with the probability of such jumps. […] measure, in some way at present unknown, the probabilities of jumps in the velocity and perhaps also in the position of an electron which forms part of an atomic system.” And in Bohr et al. [6, p.75] (1924), we read: ”Ferner nehmen wir an, dass das Vorkommen von Übergangsprozessen, sowohl für das gegebene Atom selbst wie für die anderen Atome, mit denen es kommuniziert, mit diesem Mechanismus durch Wahrscheinlichkeitsgesetze verknüpft ist, die den Gesetzen der Einsteinschen Theorie für die von äusserer Strahlung induzierten Übergänge zwischen stationären Zuständen analog sind.”
In February 2026, this interpretation was already semiquantitative, valid for large quantum numbers: ”Nach der Quantentheorie kann sich der innere Zustand eines Atoms nur sprunghaft ändern. Es sind nur die diskreten stationären Zustände der Atome möglich. An Stelle der klassischen kontinuierlichen Energieaufnahme haben wir daher eine diskontinuierliche in einzelnen Sprüngen, für die gewisse Übergangswahrscheinlichkeiten bestehen. Es müssen dann zwischen beiden korrespondenzmässige Beziehungen vorhanden sein, d.h. näherungsweise, für grosse Quantenzahlen sogar exakt, Übereinstimmung zwischen der klassischen und quantentheoretischen Rechnung herrschen.” (Nordheim [49, p.512])
@@
In June 2026, Born [8, p.822] found a highly successful statistical interpretation of the then very new wave function of Schrödinger, from which he could derive very specific physical conclusions, namely the qualitative features of collision experiments (which were later verified quantitatively). This earned Born the Nobel prize in 1954 ”for his fundamental research in quantum mechanics, especially for his statistical interpretation of the wavefunction” [48].
Wallner’s Nobel prize presentation speech for the general public comments ([48]): ”Schrödinger had not solved the problem of how it is possible to make statements about the positions and velocities of particles if one knows the wave corresponding to the particle. Born provided the solution to the problem. He found that the waves determine the probability of the measuring results. For this reason, according to Born, quantum mechanics gives only a statistical description.” This is a bit inaccurate, as Born didn’t find probabilities for positions and velocities of particles (this was done a little later by Pauli) but probabilities for state transitions of particles. We shall therefore trace the history in some detial.
@@ 1.2 The Born rule after 1970
Following Neumaier & Westra [43, Part II], we give in Section 3 precise definitions that lead to the correspondence between statistical expectations and quantum expectation claimed by the following condensed form of the Born rule.
@@ 1.3 Quantum expectation values
In quantum field theory, quantum expectation values of non-Hermitian operators are routinely used; they are already needed in the definition of so-called $N$-point functions as quantum expectation values of products of field operators. That $2$-point functions are in principle observables through linear response theory shows that the operators corresponding to observable quantities need not be Hermitian. Indeed, Dirac’s 1930 quantum mechanics textbook (Dirac [17, p.28]), which introduced the name ”observable” for operators, used this terminology for arbitrary linear operators. It is therefore interesting to note that all later editions of Dirac’s textbook and all later textbooks on quantum mechanics require the restriction to Hermitian operators possessing a real spectral resolution, i.e., in modern terminology, to self-adjoint Hermitian operators. Probably Dirac became aware of the fact that when $X$ is not normal (and in particular when $X$ is defective, hence has not even a spectral resolution), it is impossible to give the quantum value $\langle X\rangle$ of $X$ a spectral interpretation – which before 1970 was an essential ingredient of all measurement-based forms of the Born rule. Only the more recent formulations (BR-POVM), (BR-C), and (BR-Q) are again free of a spectral restriction.
@@
This form of the Born rule expresses the spirit of Heisenberg’s statement (in his 1927 paper on the uncertainty relation): ” Jeder quantentheoretischen Grösse oder Matrix lässt sich eine Zahl, die ihren ’Wert’ angibt, mit einem bestimmten wahrscheinlichen Fehler zuordnen; der wahrscheinliche Fehler hängt vom Koordinatensystem ab; für jede quantentheoretische Grösse gibt es je ein Koordinatensystem, in dem der wahrscheinliche Fehler für diese Grösse verschwindet. (Heisenberg [24, p.181f])
@@ 1.4 Discussion
von Neumann [45, p.255] derives for a quantum expectation value satisfying four (for von Neumann) plausible conditions (conditions A.–D. listed and motivated on p.250–252) the necessity of the formula (3) with a Hermitian density operator $\rho$ (his $U$) of trace 1. This is abstract mathematical reasoning, still within the core of formal (uninterpreted) quantum mechanics, independent of any relation to measurement. Note that assuming von Neumann’s condition D. (which is a restricted form of condition (D) above, forces the measurements to be projective measurements. Thus he misses the POVM generalization of the Born rule since the latter violates his condition D.
@@ 2.1 The genesis of the Born rule
The term ’Born rule’ appeared quite late, apparently first in 1934 by Bauer [5, p.302] (”la règle de Born”; cf. Ziegler [64]). Before that, during the gestation period of finding the right level of generalization and interpretation, the pioneers talked more vaguely about Born’s interpretation of quantum mechanics (or of the wave function). For example, in 1927, Jordan [30, p.811] wrote about ”Born’s Deutung der Lösung [der] Schrödingergleichung”.
Wessels [58, p.187] even goes so far as to claim that ”Neither Born nor most of his contemporaries saw in it anything that even suggests the interpretation that we now associate with his name.” Thus it is interesting to consider the genesis of the Born rule, based on the early papers of the pioneers of quantum mechanics. Mehra & Rechenberg [36] give a detailed history of Born’s statistical interpretation, and Bacciagaluppi [4] expands on the early years 1926–1927. However, they miss a number of subtle points that shed new light on the early developments.
@@
The two 1926 papers by Born [7, 8] (the first being a summary of the second) introduced the probabilistic interpretation. His 1926 formulation, ”bestimmt die Wahrscheinlichkeit dafür, daß das aus der $z$-Richtung kommende Elektron in die durch $\alpha,\beta,\gamma$ bestimmte Richtung (und mit einer Phasenänderung $\delta$) geworfen wird” (Born [7, p.865f]) when rephrased in modern terminology, is a special case of the objective scattering form (BR-OSc) of the Born rule. Note that Born didn’t have the concept of an S-matrix, first introduced in 1937 by Wheeler [60]. However, the square of the absolute value the S-matrix elements for single-particle scattering is proportional111 The S-matrix elements $S_{ik}$ are the inner products $e_{i}^{*}Se_{k}$, where the $e_{i}$ are the standard basis vectors of the matrix $S$; in Born’s case the stationary states of the particle. The relation with the S-matrix is visible more explicitly in Dirac [16], who computed in 1927 what are today called the S-matrix elements in the so-called Born approximation, to the ”Ausbeutefunktion” $\Phi_{n,m}$ mentioned in Born [8, p.824], already used (with similar notation) in Nordheim [49, p.512].
@@
De Broglie uses here exactly the same terminology as Born did in his 1927 paper [9, p.168], ”Die Teilchen sind immer von einem Wellenvorgang begleitet; diese de Broglie–Schrödingerschen Wellen hängen von den Kräften ab und bestimmen durch das Quadrat ihrer Amplitude die Wahrscheinlichkeit dafür, dass ein Teilchen in dem der Welle entsprechenden Bewegungszustand vorhanden ist. In seiner letzten Mitteilung beschäftigt sich auch Schrödinger mit dem Amplitudenquadrat der Wellen und führt dafür den Ausdruck ’Gewichtsfunktion’ ein, der sich der Terminologie der Statistik schon sehr nähert.” Thus we know that ”Bewegungszustand” stands for the energy-momentum eigenstate of the particle. Indeed, ”Bewegungszustand” (”state of motion”) was at that time a technical term now no longer in use. The 1898 book by Jordan [29, p.1f] defines: ”Mit dem Worte Kraft bezeichnet man die Ursache einer Änderung des Bewegungszustandes eines Körpers. – Als Bewegungszustand ist nicht nur eine Bewegung irgend welcher Art, sondern auch die Ruhe – als Abwesenheit jeglicher Bewegung (Geschwindigkeit = 0, vgl. S. 3) – aufzufassen. […] Jeder Körper behält den Bewegungszustand, den er in irgend einem Momente hat, nach Richtung und Geschwindigkeit unverändert bei, so lange keine äussere Kraft (ändernd) auf ihn einwirkt.” Thus Bewegungszustand is a state that persists unless changed by an interaction.
In 1925, Kramers & Heisenberg [32, p.692] use the term in the quantum context for stationary states: ”die sich auf zwei Bewegungszustände beziehen, für die die Werte der Grössen $J_{1}\cdots J_{s}$ sich um $\tau_{1}h\cdots\tau_{s}h$ unterscheiden. […] die Bewegungszustände seien durch Punkte in einer $J_{1},J_{2}$-Ebene, der Zeichenebene der Figur, dargestellt.”
@@ 2.2 Early objective, measurement-independent formulations
Born’s interpretation met with strong resistance by Schrödinger, who allegedly said – as Heisenberg [26, p.6] recalls much later – at a conference in Kopenhagen in September 1926, ”Wenn es doch bei dieser verdammten Quantenspringerei bleiben soll, dann bedauere ich, dass ich mich überhaupt mit diesem Gegenstand beschäftigthabe.”. An exchange of letters between him and Born displays the two extreme poles between which the quantum community had to find a consensus in the subsequent years.
Letter of Schrödinger to Born (November 2, 1926) (von Meyenn [37, p.329]): ”Ich habe eben Ihre letzte, ebenerhaltene Arbeit über den Adiabatensatz überflogen (also noch nicht wirklich gelesen). […] Ich habe aber doch den Eindruck, dass Sie und andere, die im Wesentlichen Ihre Ansicht teilen, zu tief im Banne derjenigen Begriffe stehen (wie stationäre Zustände, Quantensprünge usw.), die sich in den letzten zwölf Jahren Bürgerrecht in unserem Denken erworben haben, um dem Versuch, aus diesem Denkschema wieder herauszukommen, volle Gerechtigkeit widerfahren zu lassen. Als Beispiel führe ich z. B. an, dass Sie zur Kritik der Annahme: im Atom seien mehrere Eigenschwingungen gleichzeitig erregt, die Frage stellen: hat es Sinn, zu sagen, das Atom befinde sich gleichzeitig in mehreren stationären Zuständen? Ganz ferne scheint Ihnen die andere naheliegende Bemerkung zu liegen: immer und überall sonst, wo wir mit schwingungsfähigen Systemen zu tun haben, schwingen dieselben im Allgemeinen nicht in der Form einer Eigenschwingung, sondern mit einem Gemisch derselben.”
Letter of Born to Schrödinger (November 6, 1926): (von Meyenn [37, p.333f]): ”Als direkteste Evidenz der stationären Zustände haben wir die Elektronenstösse, und hier denke ich vor allem an die Stösse zweiter Art. Wir können wohl als Tatsache nehmen, dass ein Elektron beim Zusammenstoss mit einem Atom, das in irgend einem Zustande ist, immer nur einen Energiezuwachs bekommt. Wenn nun in einem Atom mehrere Eigenschwingungen gleichzeitig angeregt sein könnten, wie sollte man das verstehen? Was nützt da Ihr Gegenargument, dass das Elektron selber eine ’Wellengruppe’ sein könne? Es handelt sich doch um die Atomwellen, die diskreten Stufen entsprechen. Natürlich ist es möglich zu sagen: das in mehreren Frequenzen schwingende Atom reagiert eben in besonderer Weise so, dass es nur die einer Frequenz entsprechende Energie auf das Elektron überträgt. Aber das ist doch nur eine Ausrede. Mir scheint, dass die Bohrsche Redeweise die natürliche Beschreibung und Zusammenfassung eines grossen Tatsachenbereichs ist; darum ist es Aufgabe jeder feineren Theorie, diese Redeweise zu rechtfertigen.”
@@
Letter of Pauli to Heisenberg, Oktober 19, 1926 (Pauli [51, p.347]): ”Nun ist es so: alle Diagonalelemente der Matrizen (wenigstens von Funktionen der $p$ allein oder der $q$ allein) kann man überhaupt schon jetzt kinematisch deuten. Denn man kann ja zunächst nach der Wahrscheinlichkeit fragen, dass in einem bestimmten stationären Zustand des Systems die Koordinaten $q_{k}$ seiner Teilchen ($k=1,\ldots,f$) zwischen $q_{k}$ und $q_{k}+dq_{k}$ liegen. Die Antwort hierauf ist $|\psi(q_{1},\ldots,q_{f})|^{2}dq_{1}\ldots dq_{f}$, wenn $\psi$ die Schrödingersche Eigenfunktion ist.”
The position probability density interpretation was published in the next year in a footnote of Pauli [50, p.83,Footnote 1]: ”Wir wollen diese […] Funktion im Sinne der von Born in seiner Stoßmechanik [here he cites [7, 8]] vertretenen Auffassung des ”Gespensterfeldes” folgendermaßen deuten: Es ist $|\psi(q_{1}\ldots q_{f})|^{2}dq_{1}\ldots dq_{f}$ die Wahrscheinlichkeit dafür, daß im betreffenden Quantenzustand des Systems diese Koordinaten sich zugleich im betreffenden Volumenelement $dq_{1}\ldots dq_{f}$ des Lageraums befinden.” Apart from its objective formulation (no reference to measurement), this is a special case of the universal formulation (BR-US) of the Born rule.
The 1927 paper by von Neumann [44, p.45] generalized Pauli’s statement to arbitrary selfadjoint operators, again stated as an objective (i.e., measurement independent) interpretation. For discrete energy spectra and their energy levels, we still read p.48: ”unquantized states are impossible” (”nicht gequantelte Zustände sind unmöglich”). Thus he follows Born’s view that all physically possible states are stationary.

### v2

@@
Abstract. Details of the contents and the formulations of the Born rule changed considerably from its inception by Born in 1926 to the present day. This paper traces the early history of the Born rule 100 years ago, its generalization (essential for today’s quantum optics and quantum information theory) to POVMs around 50 years ago, and a modern derivation from an intuitive definition of the notion of a quantum detector. It is based to a large extent on little known results from the recent books ’Coherent Quantum Physics’ (2019) by A. Neumaier and ’Algebraic Quantum Physics, Vol. 1’ (2024) by A. Neumaier and D. Westra,
Also discussed is the extent to which the various forms of the Born rule have, like any other statement in physics, a restricted domain of validity, which leads to problems when applied outside this domain.
@@
This research received no funding.
Acknowledgment. I recall with pleasure heated discussions with Francois Ziegler, several years ago, about the detailed history of the Born rule.
@@ 1 Ten formulations of the Born rule
For example, we read in Jeans [35, p.91] (1924): ”Obviously quantum-dynamics must be intimately concerned with the probability of such jumps. […] measure, in some way at present unknown, the probabilities of jumps in the velocity and perhaps also in the position of an electron which forms part of an atomic system.” And in Bohr et al. [7, p.75] (1924), we read: ”Ferner nehmen wir an, dass das Vorkommen von Übergangsprozessen, sowohl für das gegebene Atom selbst wie für die anderen Atome, mit denen es kommuniziert, mit diesem Mechanismus durch Wahrscheinlichkeitsgesetze verknüpft ist, die den Gesetzen der Einsteinschen Theorie für die von äusserer Strahlung induzierten Übergänge zwischen stationären Zuständen analog sind.”
In February 1926, this interpretation was already semiquantitative, valid for large quantum numbers: ”Nach der Quantentheorie kann sich der innere Zustand eines Atoms nur sprunghaft ändern. Es sind nur die diskreten stationären Zustände der Atome möglich. An Stelle der klassischen kontinuierlichen Energieaufnahme haben wir daher eine diskontinuierliche in einzelnen Sprüngen, für die gewisse Übergangswahrscheinlichkeiten bestehen. Es müssen dann zwischen beiden korrespondenzmässige Beziehungen vorhanden sein, d.h. näherungsweise, für grosse Quantenzahlen sogar exakt, Übereinstimmung zwischen der klassischen und quantentheoretischen Rechnung herrschen.” (Nordheim [60, p.512])
@@
In June 1926, Born [9, p.822] found a highly successful statistical interpretation of the then very new wave function of Schrödinger, from which he could derive very specific physical conclusions, namely the qualitative features of collision experiments (which were later verified quantitatively). This earned Born the Nobel prize in 1954 ”for his fundamental research in quantum mechanics, especially for his statistical interpretation of the wavefunction” [59].
Wallner’s Nobel prize presentation speech for the general public comments ([59]): ”Schrödinger had not solved the problem of how it is possible to make statements about the positions and velocities of particles if one knows the wave corresponding to the particle. Born provided the solution to the problem. He found that the waves determine the probability of the measuring results. For this reason, according to Born, quantum mechanics gives only a statistical description.” This is a bit inaccurate, as Born didn’t find probabilities for positions and velocities of particles (this was done a little later by Pauli) but probabilities for state transitions of particles. This motivates a detailed study of the history of the statistical interpretation of quantum mechanics.
This paper traces the early history of the Born rule 100 years ago, its generalization (essential for today’s quantum optics and quantum information theory) to POVMs 50 years ago, and a modern derivation from an intuitive definition of the notion of a quantum detector. It is based to a large extent on little known results from my recent books (Neumaier [52] and Neumaier & Westra [54]) and three unpublished preprints (Neumaier [50, 51, 53]).
@@ 1.2 The Born rule after 1970
Following Neumaier & Westra [54, Part II], we give in Section 3 precise definitions that lead to the correspondence between statistical expectations and quantum expectation claimed by the following condensed form of the Born rule.
@@ 1.3 Quantum expectation values
In quantum field theory, quantum expectation values of non-Hermitian operators are routinely used; they are already needed in the definition of so-called $N$-point functions as quantum expectation values of products of field operators. That $2$-point functions are in principle observables through linear response theory shows that the operators corresponding to observable quantities need not be Hermitian. Indeed, Dirac’s 1930 quantum mechanics textbook (Dirac [21, p.28]), which introduced the name ”observable” for operators, used this terminology for arbitrary linear operators. It is therefore interesting to note that all later editions of Dirac’s textbook and all later textbooks on quantum mechanics require the restriction to Hermitian operators possessing a real spectral resolution, i.e., in modern terminology, to self-adjoint Hermitian operators. Probably Dirac became aware of the fact that when $X$ is not normal (and in particular when $X$ is defective, hence has not even a spectral resolution), it is impossible to give the quantum value $\langle X\rangle$ of $X$ a spectral interpretation – which before 1970 was an essential ingredient of all measurement-based forms of the Born rule. Only the more recent formulations (BR-POVM), (BR-C), and (BR-Q) are again free of a spectral restriction.
@@
This form of the Born rule expresses the spirit of Heisenberg’s statement (in his 1927 paper on the uncertainty relation): ” Jeder quantentheoretischen Grösse oder Matrix lässt sich eine Zahl, die ihren ’Wert’ angibt, mit einem bestimmten wahrscheinlichen Fehler zuordnen; der wahrscheinliche Fehler hängt vom Koordinatensystem ab; für jede quantentheoretische Grösse gibt es je ein Koordinatensystem, in dem der wahrscheinliche Fehler für diese Grösse verschwindet. (Heisenberg [31, p.181f])
@@ 1.4 Discussion
von Neumann [56, p.255] derives for a quantum expectation value satisfying four (for von Neumann) plausible conditions (conditions A.–D. listed and motivated on p.250–252) the necessity of the formula (3) with a Hermitian density operator $\rho$ (his $U$) of trace 1. This is abstract mathematical reasoning, still within the core of formal (uninterpreted) quantum mechanics, independent of any relation to measurement. Note that assuming von Neumann’s condition D. (which is a restricted form of condition (D) above, forces the measurements to be projective measurements. Thus he misses the POVM generalization of the Born rule since the latter violates his condition D.
@@ 2.1 The genesis of the Born rule
The term ’Born rule’ appeared quite late, apparently first in 1934 by Bauer [6, p.302] (”la règle de Born”; cf. Ziegler [76]). Before that, during the gestation period of finding the right level of generalization and interpretation, the pioneers talked more vaguely about Born’s interpretation of quantum mechanics (or of the wave function). For example, in 1927, Jordan [37, p.811] wrote about ”Born’s Deutung der Lösung [der] Schrödingergleichung”.
Wessels [70, p.187] even goes so far as to claim that ”Neither Born nor most of his contemporaries saw in it anything that even suggests the interpretation that we now associate with his name.” Thus it is interesting to consider the genesis of the Born rule, based on the early papers of the pioneers of quantum mechanics. Mehra & Rechenberg [45] give a detailed history of Born’s statistical interpretation, and Bacciagaluppi [5] expands on the early years 1926–1927. However, they miss a number of subtle points that shed new light on the early developments.
@@
The two 1926 papers by Born [8, 9] (the first being a summary of the second) introduced the probabilistic interpretation. His 1926 formulation, ”bestimmt die Wahrscheinlichkeit dafür, daß das aus der $z$-Richtung kommende Elektron in die durch $\alpha,\beta,\gamma$ bestimmte Richtung (und mit einer Phasenänderung $\delta$) geworfen wird” (Born [8, p.865f]) when rephrased in modern terminology, is a special case of the objective scattering form (BR-OSc) of the Born rule. Note that Born didn’t have the concept of an S-matrix, first introduced in 1937 by Wheeler [72]. However, the square of the absolute value the S-matrix elements for single-particle scattering is proportional111 The S-matrix elements $S_{ik}$ are the inner products $e_{i}^{*}Se_{k}$, where the $e_{i}$ are the standard basis vectors of the matrix $S$; in Born’s case the stationary states of the particle. The relation with the S-matrix is visible more explicitly in Dirac [20], who computed in 1927 what are today called the S-matrix elements in the so-called Born approximation, to the ”Ausbeutefunktion” $\Phi_{n,m}$ mentioned in Born [9, p.824], already used (with similar notation) in Nordheim [60, p.512].
@@
De Broglie uses here exactly the same terminology as Born did in his 1927 paper [10, p.168], ”Die Teilchen sind immer von einem Wellenvorgang begleitet; diese de Broglie–Schrödingerschen Wellen hängen von den Kräften ab und bestimmen durch das Quadrat ihrer Amplitude die Wahrscheinlichkeit dafür, dass ein Teilchen in dem der Welle entsprechenden Bewegungszustand vorhanden ist. In seiner letzten Mitteilung beschäftigt sich auch Schrödinger mit dem Amplitudenquadrat der Wellen und führt dafür den Ausdruck ’Gewichtsfunktion’ ein, der sich der Terminologie der Statistik schon sehr nähert.” Thus we know that ”Bewegungszustand” stands for the energy-momentum eigenstate of the particle. Indeed, ”Bewegungszustand” (”state of motion”) was at that time a technical term now no longer in use. The 1898 book by Jordan [36, p.1f] defines: ”Mit dem Worte Kraft bezeichnet man die Ursache einer Änderung des Bewegungszustandes eines Körpers. – Als Bewegungszustand ist nicht nur eine Bewegung irgend welcher Art, sondern auch die Ruhe – als Abwesenheit jeglicher Bewegung (Geschwindigkeit = 0, vgl. S. 3) – aufzufassen. […] Jeder Körper behält den Bewegungszustand, den er in irgend einem Momente hat, nach Richtung und Geschwindigkeit unverändert bei, so lange keine äussere Kraft (ändernd) auf ihn einwirkt.” Thus Bewegungszustand is a state that persists unless changed by an interaction.
In 1925, Kramers & Heisenberg [39, p.692] use the term in the quantum context for stationary states: ”die sich auf zwei Bewegungszustände beziehen, für die die Werte der Grössen $J_{1}\cdots J_{s}$ sich um $\tau_{1}h\cdots\tau_{s}h$ unterscheiden. […] die Bewegungszustände seien durch Punkte in einer $J_{1},J_{2}$-Ebene, der Zeichenebene der Figur, dargestellt.”
@@ 2.2 Early objective, measurement-independent formulations
Born’s interpretation met with strong resistance by Schrödinger, who allegedly said – as Heisenberg [33, p.6] recalls much later – at a conference in Kopenhagen in September 1926, ”Wenn es doch bei dieser verdammten Quantenspringerei bleiben soll, dann bedauere ich, dass ich mich überhaupt mit diesem Gegenstand beschäftigthabe.”. An exchange of letters between him and Born displays the two extreme poles between which the quantum community had to find a consensus in the subsequent years.
Letter of Schrödinger to Born (November 2, 1926) (von Meyenn [46, p.329]): ”Ich habe eben Ihre letzte, ebenerhaltene Arbeit über den Adiabatensatz überflogen (also noch nicht wirklich gelesen). […] Ich habe aber doch den Eindruck, dass Sie und andere, die im Wesentlichen Ihre Ansicht teilen, zu tief im Banne derjenigen Begriffe stehen (wie stationäre Zustände, Quantensprünge usw.), die sich in den letzten zwölf Jahren Bürgerrecht in unserem Denken erworben haben, um dem Versuch, aus diesem Denkschema wieder herauszukommen, volle Gerechtigkeit widerfahren zu lassen. Als Beispiel führe ich z. B. an, dass Sie zur Kritik der Annahme: im Atom seien mehrere Eigenschwingungen gleichzeitig erregt, die Frage stellen: hat es Sinn, zu sagen, das Atom befinde sich gleichzeitig in mehreren stationären Zuständen? Ganz ferne scheint Ihnen die andere naheliegende Bemerkung zu liegen: immer und überall sonst, wo wir mit schwingungsfähigen Systemen zu tun haben, schwingen dieselben im Allgemeinen nicht in der Form einer Eigenschwingung, sondern mit einem Gemisch derselben.”
Letter of Born to Schrödinger (November 6, 1926): (von Meyenn [46, p.333f]): ”Als direkteste Evidenz der stationären Zustände haben wir die Elektronenstösse, und hier denke ich vor allem an die Stösse zweiter Art. Wir können wohl als Tatsache nehmen, dass ein Elektron beim Zusammenstoss mit einem Atom, das in irgend einem Zustande ist, immer nur einen Energiezuwachs bekommt. Wenn nun in einem Atom mehrere Eigenschwingungen gleichzeitig angeregt sein könnten, wie sollte man das verstehen? Was nützt da Ihr Gegenargument, dass das Elektron selber eine ’Wellengruppe’ sein könne? Es handelt sich doch um die Atomwellen, die diskreten Stufen entsprechen. Natürlich ist es möglich zu sagen: das in mehreren Frequenzen schwingende Atom reagiert eben in besonderer Weise so, dass es nur die einer Frequenz entsprechende Energie auf das Elektron überträgt. Aber das ist doch nur eine Ausrede. Mir scheint, dass die Bohrsche Redeweise die natürliche Beschreibung und Zusammenfassung eines grossen Tatsachenbereichs ist; darum ist es Aufgabe jeder feineren Theorie, diese Redeweise zu rechtfertigen.”
@@
Letter of Pauli to Heisenberg, Oktober 19, 1926 (Pauli [62, p.347]): ”Nun ist es so: alle Diagonalelemente der Matrizen (wenigstens von Funktionen der $p$ allein oder der $q$ allein) kann man überhaupt schon jetzt kinematisch deuten. Denn man kann ja zunächst nach der Wahrscheinlichkeit fragen, dass in einem bestimmten stationären Zustand des Systems die Koordinaten $q_{k}$ seiner Teilchen ($k=1,\ldots,f$) zwischen $q_{k}$ und $q_{k}+dq_{k}$ liegen. Die Antwort hierauf ist $|\psi(q_{1},\ldots,q_{f})|^{2}dq_{1}\ldots dq_{f}$, wenn $\psi$ die Schrödingersche Eigenfunktion ist.”
The position probability density interpretation was published in the next year in a footnote of Pauli [61, p.83,Footnote 1]: ”Wir wollen diese […] Funktion im Sinne der von Born in seiner Stoßmechanik [here he cites [8, 9]] vertretenen Auffassung des ”Gespensterfeldes” folgendermaßen deuten: Es ist $|\psi(q_{1}\ldots q_{f})|^{2}dq_{1}\ldots dq_{f}$ die Wahrscheinlichkeit dafür, daß im betreffenden Quantenzustand des Systems diese Koordinaten sich zugleich im betreffenden Volumenelement $dq_{1}\ldots dq_{f}$ des Lageraums befinden.” Apart from its objective formulation (no reference to measurement), this is a special case of the universal formulation (BR-US) of the Born rule.
The 1927 paper by von Neumann [55, p.45] generalized Pauli’s statement to arbitrary selfadjoint operators, again stated as an objective (i.e., measurement independent) interpretation. For discrete energy spectra and their energy levels, we still read p.48: ”unquantized states are impossible” (”nicht gequantelte Zustände sind unmöglich”). Thus he follows Born’s view that all physically possible states are stationary.